OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (1, 5) and is parallel to the line y = –3x – 3?

OpenStudy (hoblos):

parallel lines have same slope so slope=-3 y-5=-3(x-1) y-5 = -3x +3 3x+y = 8

OpenStudy (anonymous):

thank you

OpenStudy (hoblos):

yw :)